OpenStudy (donnie1999):

PLEASE HELP WILL FAN AND MEDAL! Every evening Jenna empties her pockets and puts her change into a jar. At the end of the week she counts her money. One week she had 50 coins, all of them dimes and quarters. When she added them up, she had a total of $10.25. How many dimes did jenna have?

OpenStudy (donnie1999):

@phi

OpenStudy (donnie1999):

d= number of dimes and q= number of quarters.

OpenStudy (phi):

you can solve this by finding two equations ok with your "names" now use this info: she had 50 coins can you make an equation using d, q and 50 ?

OpenStudy (donnie1999):

d+q=50?

OpenStudy (phi):

yes, the number of dimes plus the number of quarters is 50 next, how much is a dime worth ? (in dollars)

OpenStudy (donnie1999):

10 cents

OpenStudy (phi):

that's in cents. how about in dollars ?

OpenStudy (donnie1999):

0.10

OpenStudy (phi):

next idea: how much is (for example) 4 dimes worth ?

OpenStudy (donnie1999):

0.40

OpenStudy (phi):

and how did you figure that out ?

OpenStudy (donnie1999):

4*10

OpenStudy (phi):

you either added up 0.10 four times: 0.10 + 0.10 + 0.10 + 0.10 or you multiplied 4 * 0.10 the 2nd way is faster

OpenStudy (phi):

now if we have "d" dimes we still use the same idea: d dimes are worth d*0.10

OpenStudy (phi):

how much are q quarters worth (in dollars, not cents) ?

OpenStudy (donnie1999):

0.25

OpenStudy (donnie1999):

so q*0.25?

OpenStudy (phi):

yes 0.25 q now we (sort of) know how much money we have: 0.10d + 0.25q they tell us how much that adds up to. can you write an equation ?

OpenStudy (donnie1999):

0.10(d)+0.25(q)=$10.25

OpenStudy (phi):

yes. now we have two equations and two unknowns (if we have an equation for each unknown, we have enough info to solve the "system") d+q=50 0.10(d)+0.25(q)=10.25 (I left off the dollar sign, but remember we are using dollars) once we get to here, it's algebra and we use everything we know about solving system of equations (if we remember!)

OpenStudy (phi):

if you multiply the bottom equation by 100, we can "get rid of" the decimals: that means multiply both sides and all terms by 100

OpenStudy (phi):

in other words, the 2nd equation can also be 10 d + 25 q = 1025 I notice each "coefficient" can be divided by 5, so I would do that i.e. divide both sides and all terms by 5. we get 2 d + 5 q = 205 now we have 1 d + 1 q = 50 2 d + 5 q =205

OpenStudy (donnie1999):

ok so i just solved it and got d=15 q=35?

OpenStudy (phi):

yes, that looks good. and they want the number of dimes

OpenStudy (donnie1999):

so its 15 thanks:D